A power series solution about x=0 of the differential equation y"-y=0

answers

3 H, - A EWE BW H22 1 - 2 22 ปอ: 388 1 SWPWE PW PL SWE 2, 2 2พะ 1) 9 - 20 4
Answer #1

Solving the power series solution we getThe given differential equation is ylyzo comparing with bolny+ Pilnyt P (n)y=0 we get Polny = 1 => Poln) to at nao >n=0 isEquating the co-efficent of n we get, 693 -c=0 693 =ą az G Equating the co-efficent of an we get (n+1)(n+2) antz - an so an aaz as (6 (5) az als an * a5 1 20 [ azzal y as 120 put n=4 in (iii) we get an (4+1) (472) a4+2 Ox4 ( 516) au > a= 30 ao co au22n > y = 20 5 2nti su ta (29 2n+1)! n=0 n=c .: The 3rd option is correct.